OpenStudy (anonymous):

write the expression as a sum/differerence of logs, with variables to the first degree

OpenStudy (anonymous):

\[2\ln6-1/2\ln36\]

OpenStudy (anonymous):

ln 6^2 - ln 36^(1/2) = ln 36 - ln 6 = ln (36/6) = ln 6